# Tensor Product Attention Is All You Need

### v5

@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language, ; chowdhery2022palm, ; touvron2023llama, ; bubeck2023sparks, ). As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion. However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o, ; liu2024kivi, ). Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating, ) or token eviction strategies (zhang2023h2o, ; xiao2023efficient, ; ribar2023sparq, ), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (he2024fastdecode, ), at the expense of increased I/O latency. Attention variants like Multi-Query Attention (MQA) (shazeer2019fast, ) and Grouped-Query Attention (GQA) (ainslie2023gqa, ) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora, ) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime. The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek, ) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer, ) integration, necessitating additional position-encoded parameters per head.
To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1. TPA is a novel attention mechanism that employs tensor factorizations for queries (Q), keys (K), and values (V). By dynamically factorizing activations rather than static weights (as in LoRA), TPA constructs low-rank, contextual representations. This approach substantially reduces KV cache memory usage while offering improved representational capacity. In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance.
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ), Qwen (bai2023qwen, ), and Gemma (team2024gemma, ).
@@
    We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions. This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention, ), MHA, MQA, GQA, and MLA, while also improving performance. In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
@@
    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer, ), facilitating its easy adoption in popular foundation model architectures like LLaMA and Gemma.
@@ 2 Background
In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ), and introduce key notations. Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in the Appendix F.
@@ 2.1 Scaled Dot-Product Attention
Scaled dot-product attention (vaswani2017attention, ) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$). It produces a weighted combination of the values ($\mathbf{V}$). Formally, the attention output is:
@@ 2.2 Multi-Head Attention (MHA)
Multi-Head Attention (MHA) (vaswani2017attention, ) extends scaled dot-product attention by dividing the model’s internal representation into several heads. Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces. For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
Parameter Initialization. We use Xavier initialization (glorot2010understanding, ) for the factor weight matrices; details are in the Appendix G.
@@ 4.3 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama, ) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu, ; touvron2023llama, ).
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (su2024roformer, ) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
@@ 6.1 Language Modeling Tasks
All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022, ), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu, ). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (touvron2023llama, ) with SwiGLU activation (shazeer2020glu, ) and RoPE embeddings (su2024roformer, ), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention ) with Multi-Query Attention (MQA; shazeer2019fast ), Grouped Query Attention (GQA; ainslie2023gqa ), or Multi-head Latent Attention (MLA; liu2024deepseek ). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). We pretrain all models with 50B tokens using 1 epoch. Details on architecture hyperparameters and training hardware are shown in Appendix H.1.
Training & Validation Curves. Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Training loss curves are provided in Appendix Figure 3. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix G) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses. For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek, ) (blue curves) generally trains more slowly and yields higher validation losses.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC, ), BoolQ (BoolQ, ), HellaSwag (hellaswag, ), OBQA (OpenBookQA, ), PIQA (PIQA, ), WinoGrande (WinoGrande, ), and MMLU (MMLU, ), using the lm-evaluation-harness codebase (eval-harness, ). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 11 and 12 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 131415 and 16 in the appendix.
@@ 6.2 Experimental Results on FlashTPA Decoding
This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms. We benchmark FlashTPA against FlashMHA (shah2024flashattention, ), FlashGQA, FlashMQA, and FlashMLA (flashmla2025, ). It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton, ). While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential. Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
@@ B.1 Triton FlashTPA Decoding Kernel
We implement the experiments using Triton [57]; Algorithm 3 sketches the kernel corresponding to Algorithm 2. The provided kernel outline specializes to the frequently used case $R_{K}{=}R_{V}{=}1$; general ranks follow by tiling over $R_{K},R_{V}$ and replacing the rank‑1 vector–matrix products with the corresponding small GEMMs in steps $S1/S2$ and the value mixing path.
@@ Appendix E More Related Work
Transformers and Attention. As a sequence-to-sequence architecture, Transformer [60] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [10, 49, 16, 30, 27, 31], kernel-based projections [11], and linearized transformers [59, 25, 44, 69, 54, 67]. To decrease memory usage and circumvent the limitation of memory bandwidth in training, [46] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [2] divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 [32] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. Concurrently, [21] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q. Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
@@
KV Cache Optimization. During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [40]. This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [1]. Consequently, various studies have explored methods to mitigate these issues. These include KV cache eviction strategies that discard less significant tokens [70, 62, 8, 1], dynamic sparse attention mechanisms focusing on selected keys and values [42, 55, 50], offloading the KV cache to CPU memory [17, 26, 53], and quantizing the KV cache [61, 34, 19]. In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
@@
Low-Rank Factorizations. Low-rank approximations are widely used to compress model parameters and reduce computational complexity. Notable examples include LoRA [20], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [28], MoRA [22]), long-context training (LongLoRA [9], SinkLoRA [66]), and continual training (InfLoRA [29], GS-LoRA [71], I-LoRA [41]). These methods generally produce static low-rank expansions that are independent of the input context. Theoretical justifications for the expressiveness of low-rank approximations have been provided by [38, 65]. Initialization strategies for these factorization matrices have also been explored: OLoRA [7] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [48] employs LDU-decomposition to accelerate LoRA training. Furthermore, AdaLoRA [68] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks. TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.

### v6

@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language; chowdhery2022palm; touvron2023llama; bubeck2023sparks). As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion. However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o; liu2024kivi). Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating) or token eviction strategies (zhang2023h2o; xiao2023efficient; ribar2023sparq), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (he2024fastdecode), at the expense of increased I/O latency. Attention variants like Multi-Query Attention (MQA) (shazeer2019fast) and Grouped-Query Attention (GQA) (ainslie2023gqa) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime. The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer) integration, necessitating additional position-encoded parameters per head.
To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1. TPA is a novel attention mechanism that employs tensor factorizations for queries (Q), keys (K), and values (V). By dynamically factorizing activations rather than static weights (as in LoRA), TPA constructs low-rank, contextual representations. This approach substantially reduces KV cache memory usage while offering improved representational capacity. In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance. A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer) and any possible position encodings, enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama), Qwen (bai2023qwen), and Gemma (team2024gemma).
@@
    We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions. This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention), MHA, MQA, GQA, and MLA, while also improving performance. In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
@@
    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer) and any possible position encodings as well as output gate and KV shifting, facilitating its easy adoption in popular foundation model architectures like LLaMA, Gemma, and Qwen.
@@ 2 Background
In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention), and introduce key notations. Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast), Grouped Query Attention (GQA) (ainslie2023gqa), Multi-head Latent Attention (MLA) (liu2024deepseek; liu2024deepseekv3), and Rotary Position Embedding (RoPE) (su2024roformer) are further discussed in the Appendix F.
@@ 2.1 Scaled Dot-Product Attention
Scaled dot-product attention (vaswani2017attention) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$). It produces a weighted combination of the values ($\mathbf{V}$). Formally, the attention output is:
@@ 2.2 Multi-Head Attention (MHA)
Multi-Head Attention (MHA) (vaswani2017attention) extends scaled dot-product attention by dividing the model’s internal representation into several heads. Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces. For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
Parameter Initialization. We use Xavier initialization (glorot2010understanding) for the factor weight matrices; details are in the Appendix G.
@@ 4.3 Model Architectures
We propose a new architecture called Tensor ProducT ATTenTion Transformer (T6), which uses our Tensor Product Attention (TPA) in place of standard MHA (multi-head attention) or GQA (grouped-query attention). Building upon the query, key, and value tensors $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ defined in Section 3.1, T6 utilizes the overall architecture of LLaMA (touvron2023llama) while changing the self-attention block to our TPA-based version. The feed-forward network (FFN) adopts a SwiGLU layer, as in (shazeer2020glu; touvron2023llama).
Rotary Positional Embedding (RoPE). As discussed in Section 3.2, RoPE (su2024roformer) is applied to the $\mathbf{Q}$ and $\mathbf{K}$. Within TPA, we pre-rotate the factor $\mathbf{b}^{Q}_{t}(\mathbf{x}_{t})$ and $\mathbf{b}^{K}_{s}(\mathbf{x}_{s})$ directly, so that each $\mathbf{K}_{s}$ is already rotated prior to caching, see Equation (3.5) and Theorem 3.1.
@@ 6.1 Language Modeling Tasks
All experiments reported in this paper are implemented based on the nanoGPT codebase (Karpathy2022), and we pretrain our models using the FineWeb-Edu 100B dataset (lozhkov2024fineweb-edu). The dataset contains 100 billion tokens for training and 0.1 billion tokens for validation. We compare T6 against the baseline Llama architecture (touvron2023llama) with SwiGLU activation (shazeer2020glu) and RoPE embeddings (su2024roformer), as well as Llama variants that replace Multi-Head Attention (MHA; vaswani2017attention) with Multi-Query Attention (MQA; shazeer2019fast), Grouped Query Attention (GQA; ainslie2023gqa), or Multi-head Latent Attention (MLA; liu2024deepseek). In our experiments, the number of heads $h$ is adjusted for each attention mechanism to ensure that all attention mechanisms have the same number of parameters as the standard Multi-Head Attention (MHA), which has $4d_{\text{model}}^{2}$ parameters per attention layer. We train models at four scales: small (124M parameters), medium (353M), large (773M), and XL (1.5B). We pretrain all models with 50B tokens using 1 epoch. Details on architecture hyperparameters and training hardware are shown in Appendix H.1.
Training & Validation Curves. Figure 4 compares validation loss curves for the medium (353M), large (773M), and XL (1.5B) models on FineWeb-Edu-100B. Training loss curves are provided in Appendix Figure 3. Overall, TPA (red curves) and its simpler variant TPA-KVonly (pink curves) (see Appendix G) converge as fast as or faster than the baselines (MHA, MQA, GQA, MLA) while also achieving visibly lower final validation losses. For instance, in Figure 4(b), TPA and TPA-KVonly remain below the MHA baseline in terms of validation loss at nearly all training stages. Meanwhile, Multi-Head Latent Attention (MLA) (liu2024deepseek) (blue curves) generally trains more slowly and yields higher validation losses.
@@
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (ARC), BoolQ (BoolQ), HellaSwag (hellaswag), OBQA (OpenBookQA), PIQA (PIQA), WinoGrande (WinoGrande), and MMLU (MMLU), using the lm-evaluation-harness codebase (eval-harness). For ARC-E, ARC-C, HellaSwag, OBQA, PIQA, and SciQ, we report accuracy norm; for other tasks, we report standard accuracy. Due to the page limitation, we only display the zero-shot evaluation results of medium and large models here in Tables 2 and 3. Zero-shot evaluation of small and XL models are displayed in Tables 11 and 12 in the appendix. Moreover, we also present 2-shot evaluation results in Tables 131415 and 16 in the appendix.
@@ 6.2 Experimental Results on FlashTPA Decoding
This section presents an evaluation of FlashTPA’s decoding time in comparison to several other optimized attention mechanisms. We benchmark FlashTPA against FlashMHA (shah2024flashattention), FlashGQA, FlashMQA, and FlashMLA (flashmla2025). It is important to note that our current FlashTPA implementation utilizes Triton (tillet2019triton). While the compared methods are typically available as highly optimized CUDA kernels, these experiments provide initial insights into FlashTPA’s potential. Development of a CUDA-based FlashTPA kernel is ongoing and is expected to yield further performance improvements.
@@ B.1 Triton FlashTPA Decoding Kernel
We implement the experiments using Triton [tillet2019triton]; Algorithm 3 sketches the kernel corresponding to Algorithm 2. The provided kernel outline specializes to the frequently used case $R_{K}{=}R_{V}{=}1$; general ranks follow by tiling over $R_{K},R_{V}$ and replacing the rank‑1 vector–matrix products with the corresponding small GEMMs in steps $S1/S2$ and the value mixing path.
@@ Appendix E More Related Work
Transformers and Attention. As a sequence-to-sequence architecture, Transformer [vaswani2017attention] introduced Multi-Head Attention (MHA), enabling more effective capture of long-range dependencies. Subsequent work has explored a variety of attention mechanisms aimed at improving scalability and efficiency, including sparse patterns [child2019generating, DBLP:conf/iclr/Shi0LR0LJ23, DBLP:conf/iclr/HanJKMWZ24, liang2024conv, li2024tighter, liang2024beyond], kernel-based projections [choromanski2020rethinking], and linearized transformers [tsai2019transformer, katharopoulos2020transformers, schlag2021linear, zhang2023trained, sun2023retentive, DBLP:conf/iclr/ZhangBKR24]. To decrease memory usage and circumvent the limitation of memory bandwidth in training, [shazeer2019fast] proposed Multi-Query Attention (MQA) where multiple query heads share the same key head and value head. To tackle the issue of quality degradation and instability in training, Grouped-Query Attention (GQA) [ainslie2023gqa] divides queries into several groups, and each group of queries shares a single key head and value head. Recently, DeepSeek-V2 [liu2024deepseek] applied multihead latent attention (MLA) to achieve better performance than MHA while reducing KV cache in inference time by sharing the same low-rank representation of key and value. Concurrently, [hu2024multi] proposed Multi-matrix Factorization Attention (MFA), which can be simply seen as MQA with low-rank factorized Q. Compared to the approaches above, TPA applied contextual tensor decompositions to represent queries, keys, and values activations compactly, achieving better reduction on the size of KV cache with improved performance.
@@
KV Cache Optimization. During the auto-regressive inference of Transformers, key and value (KV) tensors from previous tokens are cached to avoid recomputation, a technique first proposed by [ott2019fairseq]. This Key-Value (KV) cache, while crucial for efficiency, consumes significant memory and can introduce latency bottlenecks due to memory bandwidth limitations [adnan2024keyformer]. Consequently, various studies have explored methods to mitigate these issues. These include KV cache eviction strategies that discard less significant tokens [zhang2023h2o, xiao2023efficient, cai2024pyramidkv, adnan2024keyformer], dynamic sparse attention mechanisms focusing on selected keys and values [ribar2023sparq, tang2024quest, singhania2024loki], offloading the KV cache to CPU memory [he2024fastdecode, lee2024infinigen, sun2024shadowkv], and quantizing the KV cache [xiao2023smoothquant, liu2024kivi, hooper2024kvquant]. In contrast to these approaches, TPA focuses on reducing the intrinsic size of the KV cache by employing tensor-decomposed key and value representations.
@@
Low-Rank Factorizations. Low-rank approximations are widely used to compress model parameters and reduce computational complexity. Notable examples include LoRA [hu2021lora], which factorizes weight updates during fine-tuning, and its derivatives tailored for various training scenarios such as efficient pretraining (ReLoRA [lialin2023relora], MoRA [jiang2024mora]), long-context training (LongLoRA [chen2023longlora], SinkLoRA [zhang2024sinklora]), and continual training (InfLoRA [liang2024inflora], GS-LoRA [zhao2024continual], I-LoRA [ren2024analyzing]). These methods generally produce static low-rank expansions that are independent of the input context. Theoretical justifications for the expressiveness of low-rank approximations have been provided by [malladi2023kernel, zeng2023expressive]. Initialization strategies for these factorization matrices have also been explored: OLoRA [buyukakyuz2024olora] utilizes QR-decomposition of pretrained weights for improved language model performance, while LoLDU [shi2024loldu] employs LDU-decomposition to accelerate LoRA training. Furthermore, AdaLoRA [zhang2023adalora] uses Singular Value Decomposition (SVD) on pretrained weights and introduces parameter importance scores to dynamically adjust ranks. TPA, in contrast, constructs Q, K, and V tensors using contextually-aware factorizations, allowing for dynamic adaptation based on the input.